OpenStudy (blaque678):

log x is the exponent to which the base 10 must be raised to get x.So we can complete the following table for log x. 10 1/2?

OpenStudy (anonymous):

are you supposed to find \(\log(10^{\frac{1}{2}})\) ? is that the question? it is not real very clear

OpenStudy (blaque678):

yes I have the other log x but that's the only one im not sure of

OpenStudy (anonymous):

\(\log_{10}(10^x)=x\)

OpenStudy (anonymous):

more generally \(\log_b(b^x)=x\)

OpenStudy (anonymous):

that is what this line is telling you log x is the exponent to which the base 10 must be raised to get x

OpenStudy (anonymous):

since evidently you have to raise ten to the power of one half to get ten to the power of one half (kind of like who's buried in grant's tomb) then \[\log_{10}(10^{\frac{1}{2}})=\frac{1}{2}\]

OpenStudy (anonymous):

i have no idea what " 10 = 10.5" means

OpenStudy (blaque678):

I typed the wrong thing